OpenStudy (hnamowicz):

7⋅4+6−12÷47⋅4+6−12÷4

OpenStudy (hnamowicz):

31?

OpenStudy (mathstudent55):

No, You have only multiplications, divisions, additions, and subtractions. You must follow the correct order of operations. First, do multiplications and divisions in the order they appear from left to right. Then do additions and subtractions in the order they appear from left to right.

OpenStudy (hnamowicz):

ok

OpenStudy (hnamowicz):

I still got the same answer

OpenStudy (mathstudent55):

Just do the operations shown in red below. Write their results on the next line, and all numbers you haven't worked with yet, just copy them down. \(\color{red}{7 \cdot 4} + 6 - \color{red}{12\div 47} \cdot 4 + 6 - \color{red}{12 \div 4 }=\)

OpenStudy (mathstudent55):

You have a division by 47. How can you possibly end up with a whole number like 31 if you need to divide by 47?

OpenStudy (mathstudent55):

Is there an operation between the 4 and the 7 that is missing, and it is not supposed to be the number 47?